# Learning nonnegative matrix factorizations from compressed data

### v1

@@ Abstract.
Princeton University, Department of Operations Research and Financial Engineering. Emails: azc@alumni.princeton.edu, elre@princeton.edu. The authors acknowledge partial support by NSF DMS-2309685 and DMS-2108479.
@@ 1. Introduction
Low-rank approximations are arguably the main tool for simplifying and interpreting large, complex datasets. Methods based on singular value decomposition of the data matrix deliver deterministic, useful results via polynomial-time algorithms. However, for nonnegative data, spatial localization and interpretability of the features can be boosted by additionally making the factors element-wise nonnegative [23]. In the standard form, given a nonnegative matrix ${\mathbf{X}}\in{\mathbb{R}}^{m\times n}$, and a target rank $r$, the Nonnegative Matrix Factoriration (NMF) problem is the task of finding matrices ${\mathbf{U}}\in{\mathbb{R}}^{m\times r}$ and ${\mathbf{V}}\in{\mathbb{R}}^{n\times r}$ solving the problem
$\min_{{\mathbf{U}},{\mathbf{V}}\geq 0}\|{\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}.$ (1.1)
@@
The problem (1.1) is in general NP-hard [38] to solve and it never possesses a unique solution. Despite these challenges, several iterative algorithms have been developed to solve the NMF problem, including multiplicative updates (MU), alternating least squares (ALS), alternating nonnegative least squares (ANLS), and hierarchical alternating least squares (HALS). See [11, Section 3] for a discussion of these methods.
When the size of the data matrix ${\mathbf{X}}$ is large, it is challenging or impossible to store it, and even harder to use it within the optimization algorithms searching for the NMF decomposition. Note that the resulting factors ${\mathbf{U}}$ and ${\mathbf{V}}$ collectively contain only $r(n+m)$ entries which is much less than the total of $nm$ entries in ${\mathbf{X}}$ if $r\ll\min\{n,m\}$. Thus, it can be possible to store and process the resulting factors, even when processing the whole data is challenging.
@@
In this work, we develop the theory showing why and how these data-adapted sketches are useful in finding nonnegative low-rank components; and we also consider the cases when random, independent from the data, sketches can be used.
One way to sketch a structured object – in our case, a matrix – is to vectorize it and use random linear map on the resulting vector. This includes standard compressive sensing approaches for low-rank matrix recovery such as in [3, 7]. Another way (which is the focus of this work) is to consider sketches that take of the form of left and right matrix products with ${\mathbf{X}}$, e.g., ${\mathbf{A}}{\mathbf{X}}$ or ${\mathbf{X}}{\mathbf{B}}$ for appropriately sized matrices ${\mathbf{A}}$ and ${\mathbf{B}}$. Column and row sketching have been used successfully for matrix factorization and approximation problems [37, 8, 44], and its higher order analogue modewise sketching was used to speed up tensor low-rank fitting [19, 15]. An advantage of this approach is in compact sketching matrices ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ and ${\mathbf{B}}\in{\mathbb{R}}^{n\times k}$ that contain $k(n+m)$ elements together, compared to the $kmn$ entries in a linear sketch that is applied to a vectorization of the matrix ${\mathbf{X}}$ in ${\mathbb{R}}^{mn}$. Another advantage is in preserving the matrix structure of the data throughout the sketching, which is crucial in this work for integrating the compressed matrices within learning algorithms, such as the multiplicative updates algorithm.
@@ 1.1. Contributions and outline
The idea to make NMF problem scalable through randomized sketching was considered earlier. In Section 2, we review relevant related work. What was missing in the prior works is the crucial link between the algorithmic outcomes of compressed problems and their fitness for the original problem. Establishing such connection is challenging partially due to the limited theoretical guarantees for the convergence of NMF algorithms (which is essentially inevitable dues to the NP-hardness of the problem). We approach this challenge in the following way: (1) we define the compressed problems such that we can provably compare their optimal solutions with the optimal solutions to the uncompressed problem, and (2) we propose efficient algorithms for these compressed problems. Due to (1), this also means getting good solutions for the original problem.
@@
The proposed compressed problem with row and column sketching is as follows
@@
[…]
Theorem 3.1 guarantees that if ${\mathbf{X}}$ has an exact nonnegative factorization of rank $r$, then the solution to the problem above is also exact ${\mathbf{X}}=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}$ as long as the sketching dimension is at least $r$. Crucially, the matrices ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ can be generic. We explain how to solve this problem in the sketched space despite the inclusion of the regularizer terms involving orthogonal projections. Empirically, as shown in Section 5, this problem can be employed in a simplified form with $\lambda_{1}=\lambda_{2}=0$, and it is suitable for the data with approximately low nonnegative rank: if the sketching matrices are generic (in particular, not designed to approximate the range of the data), the two-sided method should be employed for tight recovery.
@@
So, what is required from the sketching matrices to work successfully within one-sided compression, (II) or (III)? Theorem 3.4 is stated for the sketching matrices with orthonormal rows: in this case, the regularizer in the form $P^{\perp}_{\mathbf{A}}{\mathbf{U}}{\mathbf{V}}^{T}=({\mathbf{I}}-{\mathbf{A}}{\mathbf{A}}^{T}){\mathbf{U}}{\mathbf{V}}^{T}$ can be conveniently incorporated in the efficient solvers. Otherwise we can use the simplified regularizer without projection operator, Theorem 3.10, where the resulting loss depends on ${\mathbf{A}}$ being approximately orthogonal (to the extent easily satisfied by generic random matrices, as described in Remark 3.11). We note that in the one-sided cases, nonzero regularization $\lambda$ is crucial both theoretically an empirically (Figure 5).
Informally, both Theorems 3.4 and 3.10 state that when we find $\tilde{{\mathbf{U}}}$ and $\tilde{{\mathbf{V}}}^{T}$ solving the compressed problems stated above, the error $\|{\mathbf{X}}-\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}\|^{2}_{F}$ depends on (a) how well an existent (unknown) solution of rank $r$ fits the uncompressed problem $\|{\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}$, (b) how well the sketching matrix approximates the range of the data $\|P^{\perp}_{\mathbf{A}}{\mathbf{X}}\|^{2}_{F}$, and (c) how close to orthonormal are the rows of ${\mathbf{A}}$. In particular, this explains and quantfies how orthogonal data-adapted measurements (e.g., constructed via the randomized rangefinder algorithm [14]) are useful in finding nonnegative low-rank components. By Corollary 3.6, in this case, the solution of the compressed problem above is exact for the original ${\mathbf{X}}$ that admits exact rank $r$ decomposition with the sketch size $k$ slightly oversamples $r$. Compared to that, data-oblivious one-sided measurements incur additional loss, both theoretically and empirically, however they can be are advantageous for other reasons. For example, they can be re-generated when needed without any access to data and they do not require an additional pass over the data to form them.
In Section 4, we move from comparing the optimal values to solving the three compressed problems above. We formulate new variants of the multiplicative updates (MU) algorithm for each of them and show that the losses are nonincreasing during their iterations in Corollaries 4.3, 4.4, and 4.5 respectively. These corollaries follow from a more general result Theorem 4.1 formulated for a generic loss function with sketching. We also briefly discuss using projected gradient descent method on our compressed objectives.
@@ 2. Related work on scalable NMF
First, several works propose iterative random sketching approach, which requires sampling new random sketching matrices as the algorithm progresses. Such works include the methods involving random projection streams [45, 46] that allow for Johnson-Lindenstrauss-type sketching matrices (random and oblivious to the data) but require multiple sketches and passes over the initial data. Similarly, in [28, 27] compressed and distributed versions of different descent methods for NMF use different compression matrices in iterations as opposed to a singular compression. Our focus is on the setting that requires one or two passes over the original data in the preprocessing stage while the algorithm that searches for nonnegative factors works solely on the compressed data.
@@
2.1. Notations
[…]
@@ 3. Compressed problems with reliable solutions
We formulate optimization problems analogous to (1.1), which do not require storing the entire data matrix ${\mathbf{X}}$ and instead use sketched measurements. This is achieved by the use of carefully chosen regularization terms. In this section, we prove that the formulated problems are guaranteed to approximately solve the original NMF problem. In the next section, we show that the standard NMF solvers are easily extendable to these new regularized objective functions.
@@ 3.1. Two-sided compression
Theorem 3.1.
Suppose ${\mathbf{X}}$ has an exact nonnegative factorization ${\mathbf{X}}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$, where ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{n\times r}$ and they are both full-rank, $r\leq\min\{n,m\}$. Let ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ be matrices of sizes $r\times m$ and $n\times r$ respectively such that ${\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}$ is invertible [^1] Note that this condition holds generically, i.e. for all but a (Lebesgue) measure-zero set of matrices.. For any $\lambda_{1},\lambda_{2}>0$, let
$\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}=\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\|{\mathbf{A}}_{1}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|_{F}^{2}$ $+\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}){\mathbf{A}}_{2}\|_{F}^{2}$ (3.1)
$+\lambda_{1}\|P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}+\lambda_{2}\|{\mathbf{U}}{\mathbf{V}}^{T}P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}\|_{F}^{2},$
@@
Remark 3.2.
We can similarly take the sketching matrices ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ of any sizes are $k_{1}\times n$ and $m\times k_{2}$ respectively with $k_{1},k_{2}\geq r$.
Remark 3.3 (Implementation considerations).
This and further objective functions are formulated to allow for memory-efficient computations. For example, in the above objective function (3.1), one need not store ${\mathbf{X}}$ and can store ${\mathbf{A}}_{1}{\mathbf{X}}$ and ${\mathbf{X}}{\mathbf{A}}_{2}$ instead. Likewise, one does not need to store or compute $P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}^{\perp}$ which is an $m\times m$ matrix, since one can instead compute
$\|P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}$ $=\textrm{Tr}({\mathbf{V}}{\mathbf{U}}^{T}P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T})$
@@
where ${\mathbf{Q}}_{2}$ is an $m\times r$ matrix with columns that form the orthonormal basis of the columns of ${\mathbf{X}}{\mathbf{A}}_{2}$, and so $P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}={\mathbf{Q}}_{2}{\mathbf{Q}}_{2}^{T}$. One can do a similar trick to compute $P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}$ in terms of an analogous matrix ${\mathbf{Q}}_{1}$ [^2] The efficient ways to find orthogonal bases of the column/row span are well-known, e.g., see the discussion in [14] Section 4.1.. Thus, one can compute the objective function of (3.1) with total storage cost $3r(n+m)$, by storing the matrices ${\mathbf{U}},{\mathbf{V}},{\mathbf{A}}_{1}{\mathbf{X}},{\mathbf{X}}{\mathbf{A}}_{2},{\mathbf{Q}}_{1},{\mathbf{Q}}_{2}$. This and similar considerations are crucial in Section 4, when we study computationally efficient iterative algorithms that solve the optimization problems on compressed data defined here.
The proof of Theorem 3.1 is loosely inspired by the row-column matrix sensing argument from [8].
@@
First, we show that the matrices ${\mathbf{U}}_{0},{\mathbf{V}}_{0}$ such that ${\mathbf{X}}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$ are not only feasible for the problem (3.1) but also give zero objective value. Indeed, since ${\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}={\mathbf{A}}_{1}{\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}{\mathbf{A}}_{2}$ is invertible, if follows that the square matrices ${\mathbf{A}}_{1}{\mathbf{U}}_{0}$ and ${\mathbf{V}}_{0}^{T}{\mathbf{A}}_{2}$ are also invertible. So, from ${\mathbf{X}}{\mathbf{A}}_{2}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}{\mathbf{A}}_{2}$ we can then compute ${\mathbf{U}}_{0}={\mathbf{X}}{\mathbf{A}}_{2}({\mathbf{V}}_{0}^{T}{\mathbf{A}}_{2})^{-1}$ and similarly ${\mathbf{V}}_{0}^{T}=({\mathbf{A}}_{1}{\mathbf{U}}_{0})^{-1}{\mathbf{A}}_{1}{\mathbf{X}}$. Then we have
${\mathbf{X}}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}={\mathbf{X}}{\mathbf{A}}_{2}({\mathbf{V}}_{0}^{T}{\mathbf{A}}_{2})^{-1}({\mathbf{A}}_{1}{\mathbf{U}}_{0})^{-1}{\mathbf{A}}_{1}{\mathbf{X}}={\mathbf{X}}{\mathbf{A}}_{2}({\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2})^{-1}{\mathbf{A}}_{1}{\mathbf{X}}=:{\mathbf{Y}}.$
This implies
${\mathbf{X}}P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}$ $={\mathbf{Y}}P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}$
$={\mathbf{X}}{\mathbf{A}}_{2}({\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2})^{-1}{\mathbf{A}}_{1}{\mathbf{X}}({\mathbf{I}}-({\mathbf{A}}_{1}{\mathbf{X}})^{T}({\mathbf{A}}_{1}{\mathbf{X}}({\mathbf{A}}_{1}{\mathbf{X}})^{T})^{-1}{\mathbf{A}}_{1}{\mathbf{X}})=\mathbf{0},$
and similarly $P_{{\mathbf{A}}_{2}^{T}{\mathbf{X}}^{T}}^{\perp}{\mathbf{X}}=\mathbf{0}$ and matrices ${\mathbf{U}}_{0},{\mathbf{V}}_{0}$ give zero objective value.
Then, since $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ are optimal for (3.1), they must also result in objective value 00 and all four summands in (3.1) vanish:
${\mathbf{A}}_{1}({\mathbf{X}}-\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T})=({\mathbf{X}}-\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T}){\mathbf{A}}_{2}=\mathbf{0},$ (3.2)

### v2

@@ 1. Introduction
Low-rank approximations are arguably the main tool for simplifying and interpreting large, complex datasets. Methods based on singular value decomposition of the data matrix deliver deterministic, useful results via polynomial-time algorithms. However, for nonnegative data, spatial localization and interpretability of the features can be boosted by additionally making the factors element-wise nonnegative [24]. In the standard form, given an element-wise nonnegative matrix ${\mathbf{X}}\in{\mathbb{R}}_{{\color[rgb]{0,0,0}+}}^{m\times n}$, and a target rank $r$, the Nonnegative Matrix Factoriration (NMF) problem is the task of finding matrices ${\mathbf{U}}\in{\mathbb{R}}^{m\times r}_{{\color[rgb]{0,0,0}+}}$ and ${\mathbf{V}}\in{\mathbb{R}}^{n\times r}_{{\color[rgb]{0,0,0}+}}$ solving the problem
(1) $\min_{{\mathbf{U}},{\mathbf{V}}\geq 0}\|{\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}.$
@@
The problem (1) is in general NP-hard [41] to solve and does not possess a unique solution. However, under additional conditions, the uniqueness up to scaling and permutation holds; see, e.g., [13, Chapter 4]. Further, several iterative algorithms have been developed to solve the NMF problem, including multiplicative updates (MU), alternating least squares (ALS), alternating nonnegative least squares (ANLS), and hierarchical alternating least squares (HALS). See [13, Chapter 8] for a discussion of these methods.
When the size of the data matrix $\mathbf{X}$} is large, it is challenging or impossible to store it, and even harder to use it within the optimization algorithms searching for the NMF decomposition. Note that the resulting factors $\mathbf{U}$} and $\mathbf{V}$} collectively contain only $r(n+m)$ entries which is much less than the total of $nm$ entries in $\mathbf{X}$} if $r\ll\min\{n,m\}$. Thus, it is reasonable to ask if one can store and process the resulting factors, even when processing the whole data is challenging.
@@
Despite recent progress on many important problems using sketching, there has been little systematic work on adapting the sketching ideas to NMF. Here, we develop a theory showing why and how the data-adapted sketches as in [15] are useful in finding nonnegative low-rank components; and we also consider the cases when random, independent from the data, sketches can be used.
One way to sketch a structured object – in our case, a matrix – is to vectorize it and use random linear map on the resulting vector. This includes standard compressive sensing approaches for low-rank matrix recovery such as in [3, 7]. Another way (which is the focus of this work) is to consider sketches that take of the form of left and right matrix products with $\mathbf{X}$}, e.g., ${\mathbf{A}}_{1}{\mathbf{X}}$ or ${\mathbf{X}}{\mathbf{A}}_{2}$ for appropriately sized matrices ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$. Column and row sketching have been used successfully for matrix factorization and approximation problems [40, 8, 47], and its higher order analogue modewise sketching was used to speed up tensor low-rank fitting [20, 16]. An advantage of this approach is in compact sketching matrices ${\mathbf{A}}_{1}\in{\mathbb{R}}^{k\times m}$ and ${\mathbf{A}}_{2}\in{\mathbb{R}}^{n\times k}$ that contain $k(n+m)$ elements together, compared to the kmnkmn entries in a linear sketch that is applied to a vectorization of the matrix $\mathbf{X}$} in ${\mathbb{R}}^{mn}$. Another advantage is in preserving the matrix structure of the data throughout the sketching, which is crucial in this work for integrating the compressed matrices within learning algorithms, such as the multiplicative updates algorithm.
@@ 1.1. Contributions and outline
The idea to make NMF problem scalable through randomized sketching was considered earlier. In Section 2, we review relevant related work. What was missing in the prior works is the crucial link between the algorithmic outcomes of compressed problems and their fitness for the original problem. Establishing such connection is challenging partially due to the limited theoretical guarantees for the convergence of NMF algorithms (which is essentially inevitable due to the NP-hardness of the problem). We approach this challenge in the following way: (1) we define the compressed problems such that we can provably compare their optimal solutions with the optimal solutions to the uncompressed problem, and (2) we propose efficient algorithms for these compressed problems. Due to (1), this also means getting good solutions for the original problem.
We note that popular and practically efficient algorithms to solve the NMF problem, such as multiplicative updates algorithm (MU) studied here, do not have strong convergence guarantees, neither on a standard NMF problem nor – similarly – on its compressed counterpart. Actually, both the algorithmic guarantees for the sketched and for the full NMF problem are developed similarly. Consecutively, the rigorous link between the compressed and full NMF is obtained in terms of the optimal values, not algorithmic outcomes.
@@
The proposed compressed problem with row and column sketching is as follows:
$\text{(I)}\quad\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}=\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}$ $\|{\mathbf{A}}_{1}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|_{F}^{2}+\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}){\mathbf{A}}_{2}\|_{F}^{2}$
@@
for some $\lambda_{1},\lambda_{2}>0$. Theorem 1 guarantees that if $\mathbf{X}$} has an exact nonnegative factorization of rank $r$, then the solution to the problem above is also exact ${\mathbf{X}}=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}$ as long as the sketching dimension is at least $r$. Crucially, the matrices ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ can be generic. We explain how to solve this problem in the sketched space despite the inclusion of the regularizer terms involving orthogonal projections. Empirically, as shown in Section 5, this problem can be employed in a simplified form with $\lambda_{1}=\lambda_{2}=0$, and it is suitable for the data with approximately low nonnegative rank: if the sketching matrices are generic (in particular, not designed to approximate the range of the data), the two-sided method should be employed for tight recovery.
@@
So, what is required from the sketching matrices to work successfully within one-sided compression, (II) or (III)? Theorem 2 is stated for the sketching matrices with orthonormal rows: in this case, the regularizer in the form $P^{\perp}_{{\mathbf{A}}^{T}}{\mathbf{U}}{\mathbf{V}}^{T}=({\mathbf{I}}-{\color[rgb]{0,0,0}{\mathbf{A}}^{T}{\mathbf{A}}}){\mathbf{U}}{\mathbf{V}}^{T}$ can be conveniently incorporated in the efficient solvers. Otherwise we can use the simplified regularizer without projection operator, Theorem 4, where the resulting loss depends on $\mathbf{A}$} being approximately orthogonal (to the extent easily satisfied by generic random matrices, as described in Remark 4). We note that in the one-sided cases, nonzero regularization $\lambda$ is crucial both theoretically and empirically (Figure 3).
Informally, both Theorems 2 and 4 state that when we find $\tilde{{\mathbf{U}}}$ and $\tilde{{\mathbf{V}}}^{T}$ solving the compressed problems stated above, the error $\|{\mathbf{X}}-\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}\|^{2}_{F}$ depends on (a) how well an existent (unknown) solution of rank $r$ fits the uncompressed problem $\|{\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}$, (b) how well the sketching matrix approximates the range of the data $\|P^{\perp}_{{\mathbf{A}}^{T}}{\mathbf{X}}\|^{2}_{F}$, and (c) how close to orthonormal are the rows of $\mathbf{A}$}. In particular, this explains and quantfies how orthogonal data-adapted measurements (e.g., constructed via the randomized rangefinder algorithm [15]) are useful in finding nonnegative low-rank components. By Corollary 1, in this case, the solution of the compressed problem above is exact for the original $\mathbf{X}$} that admits exact rank $r$ decomposition with the sketch size $k$ slightly oversamples $r$. Compared to that, data-oblivious one-sided measurements incur additional loss, both theoretically and empirically, however they can be advantageous for other reasons. For example, they can be re-generated when needed without any access to data and they do not require an additional pass over the data to form them.
In Section 4, we move from comparing the optimal values to solving the three compressed problems above. We formulate new variants of the multiplicative updates (MU) algorithm for each of them and show that the losses are nonincreasing during their iterations in Corollaries 4, 5, and 6 respectively. These corollaries follow from a more general result Theorem 5 formulated for a generic loss function with sketching. We also briefly discuss using projected gradient descent method on our compressed objectives.
@@ 2. Related work on scalable NMF
First, several works propose iterative random sketching approach, which requires sampling new random sketching matrices as the algorithm progresses. Such works include the methods involving random projection streams [48, 49] that allow for Johnson-Lindenstrauss-type sketching matrices (random and oblivious to the data) but require multiple sketches and passes over the initial data. Similarly, in [31, 30] memory-efficient versions of NMF apply iteratively updated sketching matrices to the original data, as opposed to a single compression. Our focus is on the setting that requires one or two passes over the original data in the preprocessing stage while the algorithm that searches for nonnegative factors works solely on the compressed data.
@@
2.1. Notation
Throughout the text, matrices and vectors are denoted with bold letters. We denote Frobenius matrix norm as $\|.\|_{F}$ and the spectral (operator) norm as $\|.\|$. The matrix ${\mathbf{X}}\in{\mathbb{R}}^{m\times n}_{+}$ means it is element-wise nonnegative, the same is denoted by ${\mathbf{X}}\geq 0$ when the size of the matrix is clear from the context. Element-wise positive and negative parts of vectors and matrices are denoted as $(\cdot)_{+}=\max(\cdot,0)$ and $(\cdot)_{-}=-\min(\cdot,0)$ respectively. Further, $\circ$ denotes element-wise multiplication and // denotes element-wise division. $P_{\mathbf{Z}}$ is the linear projection operator onto the column space of a tall matrix $\mathbf{Z}$}, projection to the orthogonal complement is $P_{\mathbf{Z}}^{\perp}:={\mathbf{I}}-P_{\mathbf{Z}}$.
@@ 3. Compressed problems with reliable solutions
We formulate optimization problems analogous to (1), which do not require storing the entire data matrix $\mathbf{X}$} and instead use sketched measurements. This is achieved by the use of carefully chosen regularization terms. In this section, we prove that the formulated problems are guaranteed to approximately solve the original NMF problem. In the next section, we show that the standard NMF solvers are easily extendable to these new sketched and regularized objective functions.
@@ 3.1. Two-sided compression
Theorem 1.
Suppose $\mathbf{X}$} has an exact nonnegative factorization ${\mathbf{X}}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$, where ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{n\times r}$ with $r\leq\min\{n,m\}$. Let ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ be matrices of sizes ${\color[rgb]{0,0,0}k_{1}}\times m$ and $n\times{\color[rgb]{0,0,0}k_{2}}$ respectively such that ${\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}$ has the same rank as $\mathbf{X}$} [^1] Note that this condition holds generically, i.e., for all but a (Lebesgue) measure-zero set of matrices as long as $k_{1}$ and $k_{2}$ are both at least the rank of $\mathbf{X}$}.. For any $\lambda_{1},\lambda_{2}>0$, let
(2) $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}{\color[rgb]{0,0,0}\;\in\;}\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\|{\mathbf{A}}_{1}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|_{F}^{2}$ $+\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}){\mathbf{A}}_{2}\|_{F}^{2}$
$+\lambda_{1}\|P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}+\lambda_{2}\|{\mathbf{U}}{\mathbf{V}}^{T}P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}^{\perp}\|_{F}^{2},$
@@
Remark 1 (Implementation considerations).
This and further objective functions are formulated to allow for memory-efficient computations. For example, in the objective function (2) above, it is not necessary to store the large matrix $\mathbf{X}$}; one can instead store ${\mathbf{A}}_{1}{\mathbf{X}}$ and ${\mathbf{X}}{\mathbf{A}}_{2}$. Likewise, one does not need to store or compute $P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}$ which is an $m\times m$ matrix, since one can instead compute
$\|P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}$ $=\textrm{Tr}({\mathbf{V}}{\mathbf{U}}^{T}P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T})$
$={\color[rgb]{0,0,0}\textrm{Tr}({\mathbf{V}}{\mathbf{U}}^{T}({\mathbf{I}}-P_{{\mathbf{X}}{\mathbf{A}}_{2}}){\mathbf{U}}{\mathbf{V}}^{T})}$
@@
where ${\mathbf{Q}}_{2}$ is an $m\times k_{2}$ matrix with columns that form the orthonormal basis of the columns of ${\mathbf{X}}{\mathbf{A}}_{2}$, and so $P_{{\mathbf{X}}{\mathbf{A}}_{2}}={\mathbf{Q}}_{2}{\mathbf{Q}}_{2}^{T}$. One can do a similar trick to compute $P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}^{\perp}$ in terms of an analogous matrix ${\mathbf{Q}}_{1}$ [^2] The efficient ways to find orthogonal bases of the column/row span are well-known, e.g., see the discussion in [15] Section 4.1.. Thus, one can compute the objective function of (2) with total storage cost $3r^{\prime}(n+m)$, where $r^{\prime}=\max\{r,k_{1},k_{2}\}$, by storing the matrices ${\mathbf{U}},{\mathbf{V}},{\mathbf{A}}_{1}{\mathbf{X}},{\mathbf{X}}{\mathbf{A}}_{2},{\mathbf{Q}}_{1},{\mathbf{Q}}_{2}$. This and similar considerations are crucial in Section 4, when we study computationally efficient iterative algorithms that solve the optimization problems on compressed data defined here.
The proof of Theorem 1 is loosely inspired by the row-column matrix sensing argument from [8].
@@
First, we show that the matrices ${\mathbf{U}}_{0},{\mathbf{V}}_{0}$ such that ${\mathbf{X}}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$ are not only feasible for the problem (2) but also give zero objective value. Indeed, since ${\mathbf{A}}_{1}{\mathbf{X}}{\mathbf{A}}_{2}$ has the same rank as $\mathbf{X}$}, it follows that the matrices ${\mathbf{A}}_{1}{\mathbf{X}}$ and ${\mathbf{X}}{\mathbf{A}}_{2}$ also have the same rank as $\mathbf{X}$}. This implies that the row space of ${\mathbf{A}}_{1}{\mathbf{X}}$ and the column space of ${\mathbf{X}}{\mathbf{A}}_{2}$ are the same as the row space and column space of $\mathbf{X}$}, respectively. Thus we have $P_{{\mathbf{X}}{\mathbf{A}}_{2}}{\mathbf{X}}=P_{{\mathbf{X}}}$ and $P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}{\mathbf{X}}=P_{{\mathbf{X}}^{T}}$. Then one can observe that
$\|{\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}^{\perp}\|_{F}=\|{\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}P_{{\mathbf{X}}^{T}}^{\perp}\|_{F}=\|{\mathbf{X}}P_{{\mathbf{X}}^{T}}^{\perp}\|_{F}=\|\mathbf{0}\|_{F}=0,$
and similarly $P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{X}}=\mathbf{0}$ and matrices ${\mathbf{U}}_{0},{\mathbf{V}}_{0}$ give zero objective value.
Then, since $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ are optimal for (2), they must also result in objective value 0 and all four summands in (2) vanish:
(3) ${\mathbf{A}}_{1}({\mathbf{X}}-\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T})=({\mathbf{X}}-\tilde{\mathbf{U}}\tilde{\mathbf{V}}^{T}){\mathbf{A}}_{2}=\mathbf{0},$